Two particles having charges π‘ž1 = 1.00 Γ— 10βˆ’19𝐢 and π‘ž2 = 2.00 Γ— 10βˆ’9𝐢 are separated by a distance 2.00m. At what point is the total electric field due to the two charges equal to zero?Β 


1
Expert's answer
2022-03-21T12:42:53-0400

"\vec{E_1}+\vec{E_2}=0" initial conditions

"E=\frac{kq}{r^2}" electric field


"\frac{q_1}{r_1^2}=\frac{q_2}{(l-r_1)^2}""r_1=\frac{l}{\sqrt{\frac{q_2}{q_1}}+1}"

Calculating

"r_1=\frac{2}{\sqrt{\frac{2\cdot10^{-19}}{1\cdot10^{-9}}}+1}\approx1.99997..."

Answer electric field equal to zero on distances 1.99 metres from q1